Some Pledges for Scientists and Engineers

Student Pugwash Pledge I promise to work for a better world, where science and technology are used in socially responsible ways. I will not use my education for any purpose intended to harm human beings or the environment. Throughout my career, I will consider the ethical implications of my work before I take action. While the demands placed upon me may be great, I sign this declaration because I recognize that individual responsibility is the first step on the path to peace.

Scientists’ and Engineers’ Pledge to Renounce Weapons of Mass Destruction I pledge never to participate in

• the design, development, testing, production, maintenance, targeting, or use of nuclear, biological, or chemical weapons or their means of delivery; or in

• research or engineering that I have reason to believe will be used by others to do so.

Humboldt pledge I ⟨name⟩ pledge to thoroughly investigate and take into account the social and environmental consequences of any job opportunity I consider.

Charles Schwartz’s pledge for UCB physics students The purpose of science should be the general enhancement of life and not the causing of harm to man. I affirm that I will uphold this principle, in teaching and in practice of my science, to the best of my ability and judgement. (1970)

note: Prof. Schwartz was formally reprimanded by the UCB Chancellor for requiring physics students in his classes to take this pledge.